# AgentLeak: A Full-Stack Benchmark for Privacy Leakage in Multi-Agent LLM Systems

### v1

@@ Abstract
Multi-agent Large Language Model (LLM) systems create privacy risks that current benchmarks cannot measure. When agents coordinate on tasks, sensitive data passes through inter-agent messages, shared memory, and tool arguments; pathways that output-only audits never inspect.
We introduce AgentLeak, to the best of our knowledge the first full-stack benchmark for privacy leakage covering internal channels, spanning 1,000 scenarios across healthcare, finance, legal, and corporate domains, paired with a 32-class attack taxonomy and three-tier detection pipeline. Testing GPT-4o, GPT-4o-mini, Claude 3.5 Sonnet, Mistral Large, and Llama 3.3 70B across 4,979 traces reveals that multi-agent configurations reduce per-channel output leakage (C1: 27.2% vs 43.2% in single-agent) but introduce unmonitored internal channels that raise total system exposure to 68.9% (OR-aggregated across C1, C2, C5). Internal channels account for most of this gap: inter-agent messages (C2) leak at 68.8%, compared to 27.2% on C1 (output channel). This means that output-only audits miss 41.7% of violations. Claude 3.5 Sonnet, which emphasizes safety alignment in its design, achieves the lowest leakage rates on both external (3.3%) and internal (28.1%) channels, suggesting that model-level safety training may transfer to internal channel protection. Across all five models and four domains, the pattern C2 $\geq$ C1 holds consistently, confirming that inter-agent communication is the primary vulnerability. These findings underscore the need for coordination frameworks that incorporate internal-channel privacy protections and enforce privacy controls on inter-agent communication.
@@ I Introduction
Multi-agent Large Language Model (LLM) systems [1, 2] are reshaping enterprise workflows from healthcare scheduling to financial compliance to legal discovery. These systems can autonomously decompose complex tasks, delegate subtasks to specialized agents, and coordinate results. As such, enterprises are increasingly deploying them on sensitive data: protected health information, financial records, legal documents.
@@
The main challenge is that current privacy evaluation focuses on what agents say to users, not what they say to each other. Existing benchmarks audit final outputs [9] or investigate models about privacy norms [10]. As such, in practice, we need agents that not only reason about privacy but execute handling complex inputs and completing multi-step tasks while appropriately managing sensitive information in action, including during internal coordination. For example, AgentDojo [9] tests prompt injection robustness but audits only final outputs. On the other hand, PrivacyLens [10] measures privacy norm awareness at the model level, not system behavior under real coordination. Safety benchmarks [29] measure refusal rates, not data flow.
@@
Motivated by this principle, we introduce AgentLeak a benchmark that directly evaluates how well multi-agent LLM systems can satisfy data minimization across all communication channels, not just final outputs. Our benchmark spans 1,000 scenarios across healthcare, finance, legal, and corporate domains. Unlike prior work that uses emulated environments [9] or probes LLMs directly [10], AgentLeak deploys agents in coordinator-worker topologies with instrumented data flows through seven distinct channels: final outputs, inter-agent messages, tool inputs/outputs, shared memory, logs, and artifacts.
@@ II-C Defense Mechanisms
Guardrail Systems. NVIDIA’s NeMo Guardrails [20] and Meta’s LlamaGuard [21] enforce policies at system interfaces. These systems enforce policies at system interfaces, but do not monitor internal communication pathways.
Privacy-Focused Prompts. Drawing on Constitutional AI [22], these encode privacy norms in natural language. The hope is that agents internalize them. In practice, agents often ignore prompt instructions when coordinating with each other.
@@
TABLE I: Benchmark Landscape: Identifying the Evaluation Gap.
Tool Multi-Agent Topology Output Channels Internal Channels Type
@@
NeMo Guardrails Defense
@@ III-B The Seven Leakage Channels
External Channels: Channel C1 captures final outputs presented to users the primary target of current safety mechanisms. Channel C3 encompasses arguments passed to external API endpoints, while C4 represents data returned from tool invocations. Channel C6 covers telemetry streams and system logs. Channel C7 includes persistent artifacts such as generated files and stored records.
Internal Channels: Channel C2 comprises inter-agent messages exchanged during task delegation and coordination. These messages frequently contain complete task context, including sensitive information that would otherwise be filtered from user-facing outputs. Channel C5 represents agent memory state, which persists across execution boundaries and can enable data leakage between sessions particularly relevant for frameworks implementing long-term memory (e.g., AutoGPT).
@@ III-C Privacy Leakage Formalization
Formally, let $\mathcal{V}=\{v_{1},\ldots,v_{n}\}$ be the private vault of $n$ sensitive fields we must protect. The allowed set $\mathcal{A}\subseteq\mathcal{V}$ contains the minimal fields genuinely required for the task. Let $\mathcal{C}=\{C_{1},\ldots,C_{7}\}$ denote our seven channels. We define a semantic similarity function $\text{sim}(v,c)$ measuring the usage of vault field $v$ in channel content $c$. A leakage event occurs when:
$\exists\,v_{i}\in(\mathcal{V}\setminus\mathcal{A}),\,\exists\,C_{j}\in\mathcal{C}:\text{sim}(v_{i},\text{content}(C_{j}))>\tau$ (1)
@@
This turns data minimization [14]—an abstract legal principle from GDPR [15], HIPAA [16], and Quebec’s Law 25 [17]—into a testable empirical criterion.
Terminology & Compliance Interpretation 1. Leakage vs. Exposure: We define exposure as the presence of sensitive data in a channel (C2/C5). We define leakage as unauthorized exposure. In this benchmark, all specific “vault” data is unauthorized for internal propagation under the Principle of Least Privilege. 2. Fair Comparison: When comparing Single- vs Multi-Agent, comparing only Final Output (C1) is the “fair” technical metric (measuring task output correctness). 3. System Risk: The “true” risk metric includes C2+C5, as these internal states are logged, stored, and potentially accessible to infrastructure providers or attackers (Lateral Movement), representing a regulatory liability (GDPR Art. 25).
Figure 1: Operational definitions for leakage and risk metrics used in this study.
Figure 2: AgentLeak framework-agnostic evaluation harness. The Software Development Kit (SDK) intercepts trace events across all seven channels (C1–C7), applies the three-tier detection pipeline, and produces standardized metrics regardless of the underlying multi-agent framework. This benchmark design enables consistent privacy evaluation across LangChain, CrewAI, AutoGPT, MetaGPT, and custom implementations.
Figure 3: AgentLeak’s multi-agent system architecture showing the seven leakage channels. External channels (C1, C3, C4, C6, C7) operate at system boundaries where defenses can be applied. Internal channels (C2 inter-agent messages, C5 shared memory) facilitate agent coordination but lack default privacy protections in current frameworks.
@@ IV-A Design Principles
Framework Independence. Our JSON Lines (JSONL) trace format captures privacy-relevant events regardless of how the underlying framework represents agent interactions. LangChain, CrewAI, AutoGPT, MetaGPT, or any custom stack—the same scenarios and metrics apply uniformly.
@@ IV-B Scenario Schema and Structure
The private vault lists every sensitive field that must be protected—social security numbers, diagnoses, account numbers, salaries, case details. Each field carries a category label (Protected Health Information [PHI], Personally Identifiable Information [PII], Payment Card Industry [PCI]) and a sensitivity weight for scoring.
@@
Ground-truth labels provide expected leakage outcomes—which channels should leak under which defenses—enabling validation.
@@ IV-C Domain Coverage
Our benchmark encompasses four industry verticals 250 scenarios per domain—selected for their stringent privacy requirements and documented consequences of data breaches.
@@
Finance (250 scenarios): Financial scenarios span know-your-customer and anti-money laundering workflows, portfolio management, transaction disputes, credit underwriting, and regulatory reporting. Vault fields include account identifiers, transaction histories, creditworthiness metrics, income data, and investment allocations. Financial services regularly involve data exchange among service providers, compliance teams, and regulatory entities, creating substantial regulatory complexity under frameworks including GLBA, PCI-DSS, and state privacy statutes.
Legal (250 scenarios): Legal practice scenarios address contract analysis, discovery document management, client onboarding, privilege protection, and case documentation. Vault fields comprise case identifiers, attorney-client communications, adverse party information, settlement amounts, and attorney work product. Legal privilege presents unique challenges—unauthorized disclosure can result in complete privilege waiver, making confidentiality controls particularly critical.
@@
Table II presents the complete scenario distribution across topology and threat level dimensions.
TABLE II: AgentLeak Scenario Distribution by Topology and Threat Level.
Dimension Category Count Percentage
@@
Multi-agent (Coordinator-Worker) 600 60%
@@ IV-D Data Generation and Validation
Explicit canaries: Unique markers, trivially detectable if leaked verbatim—baseline validation.
Format-valid fakes: Syntactically correct identifiers satisfying checksums (Luhn, SSN area codes)—testing format-based detection.
@@
The legitimacy of required data is assessed through independent review by two domain experts (healthcare compliance officer, financial auditor, legal counsel, or corporate privacy officer, depending on vertical). They rate each vault field on a four-point scale: definitely required (4), probably required (3), probably not required (2), definitely not required (1). Fields rated 3+ by both annotators go into the allowed set. Inter-rater reliability is measured via Cohen’s $\kappa$ [33]: Healthcare 0.82, Finance 0.79, Legal 0.81, Corporate 0.75—all indicating substantial agreement ($\kappa>0.6$) across verticals.
@@ IV-E Three-Tier Detection Pipeline
Tier 1 (Canary Matching): Regex search for unique markers inserted as explicit canaries. High confidence, zero false positives—validates that the capture infrastructure works.
@@ V Attack Taxonomy
Our 32-class taxonomy provides a standardized vocabulary for reproducible privacy research. We first motivate its necessity, then describe each attack family. Table III outlines the mapping of attack families to the 32-class taxonomy.
@@ V-A Motivation
Prior work on data extraction [23] and PII leakage targets static LLMs. Agentic systems introduce additional attack surfaces—tools, memory, and multi-agent coordination—that existing taxonomies do not cover. A systematic taxonomy enables:
@@ V-B Attack Families
TABLE III: Attack taxonomy: six families mapping to 32 classes.
@@
F3: Memory and Persistence Attacks (5 classes) target the temporal aspect of how agents retain state. Once sensitive data lands in persistent storage, an attacker can extract it long after the original task has completed.
@@
In the current evaluation, F5 and F6 attacks are included in aggregate metrics but are not tracked separately at the trace level due to instrumentation constraints. We report per-family ASR for F1 through F4 only. Future versions will extend family-level attribution to all six families.
Table IV maps attack families to adversary capability levels, showing which attacks are available at each threat level.
TABLE IV: Attack Family to Adversary Level Mapping.
@@ VI-B Detection Pipeline Validation
To mitigate evaluation leakage, judges only receive minimally necessary context: a trace segment and a masked scenario identifier. Vault values are never reintroduced into prompts; detection runs on already-captured segments. Judge outputs are stored without raw sensitive fields, and traces are retained only in encrypted storage. We report representative false positives (structured-but-benign identifiers) and false negatives (paraphrased medical facts) in Appendix A.
@@ VI-D Experimental Setup
The benchmark comprises 1,000 designed scenarios (250 per vertical: healthcare, finance, legal, corporate). Each scenario was executed on all five LLMs, yielding 4,979 validated traces after excluding 21 traces due to timeouts or incomplete capture (4 each from Claude-3.5-Sonnet, GPT-4o, Llama-3.3-70B, and Mistral-Large; 5 from GPT-4o-mini), resulting in a total of n=4,979n=4{,}979 model–scenario pairs. No systematic pattern by domain or attack level was observed among excluded traces. Table V summarizes the trace distribution.
TABLE V: Benchmark Execution Summary.
@@
Scenarios per vertical 250 GPT-4o 996 Finance 1,244
@@
All evaluations used $\text{temperature}=0.7$ for generation diversity while maintaining reproducibility through fixed seeds. Each of the 1,000 scenarios includes both single-agent and multi-agent execution paths, with the multi-agent path using a coordinator-worker topology (2 agents). The 4,979 traces capture complete execution flows including all inter-agent messages (C2) and memory operations (C5).

### v2

@@ Abstract
Multi-agent Large Language Model (LLM) systems create privacy risks that current benchmarks cannot measure. When agents coordinate on tasks, sensitive data passes through inter-agent messages, shared memory, and tool arguments, all pathways that output-only audits never inspect. We introduce AgentLeak, to the best of our knowledge the first full-stack benchmark for privacy leakage covering internal channels. It spans 1,000 scenarios across healthcare, finance, legal, and corporate domains, paired with a 32-class attack taxonomy and a three-tier detection pipeline. A factorial evaluation crossing five production LLMs (GPT-4o, GPT-4o-mini, Claude 3.5 Sonnet, Mistral Large, and Llama 3.3 70B) with all 1,000 scenarios, yielding 4,979 validated execution traces, reveals that multi-agent configurations reduce per-channel output leakage (C1: 27.2% vs 43.2% in single-agent) but introduce unmonitored internal channels that raise total system exposure to 68.9% (aggregated across C1, C2, C5). Internal channels account for most of this gap: inter-agent messages (C2) leak at 68.8%, compared to 27.2% on C1 (output channel). This means that output-only audits miss 41.7% of violations. Safety-aligned models achieve lower leakage on both external and internal channels, yet no model eliminates it. Across all five models and four domains, the pattern C2 $\geq$ C1 holds consistently, confirming that inter-agent communication is the primary vulnerability. These results establish that output-only auditing is fundamentally insufficient for multi-agent systems and that privacy controls must be extended to inter-agent communication channels.
Index Terms:
AI safety, autonomous agents, benchmark, contextual integrity, data minimization, data security, large language models, multi-agent systems, privacy leakage, security evaluation.
doi:
10.1109/ACCESS.2026.XXXXXXX
[…]
=-15pt
@@ I Introduction
Multi-agent Large Language Model (LLM) systems [2, 1] are reshaping enterprise workflows from healthcare scheduling to financial compliance to legal discovery. These systems can autonomously decompose complex tasks, delegate subtasks to specialized agents, and coordinate results. As such, enterprises are increasingly deploying them on sensitive data: protected health information, financial records, legal documents [3].
@@
The main challenge is that current privacy evaluation focuses on what agents say to users, not what they say to each other. Existing benchmarks audit final outputs [4] or investigate models about privacy norms [5]. As such, in practice, we need agents that not only reason about privacy but also handle complex inputs and complete multi-step tasks while appropriately managing sensitive information in action, including during internal coordination. For example, AgentDojo [4] tests prompt injection robustness but audits only final outputs. On the other hand, PrivacyLens [5] measures privacy norm awareness at the model level, not system behavior under real coordination. Safety benchmarks [7, 6] measure refusal rates, not data flow.
@@
Motivated by this principle, we introduce AgentLeak, a benchmark that directly evaluates how well multi-agent LLM systems can satisfy data minimization across all communication channels, not just final outputs. Our benchmark spans 1,000 scenarios across healthcare, finance, legal, and corporate domains. Unlike prior work that uses emulated environments [4] or probes LLMs directly [5], AgentLeak deploys agents in coordinator-worker topologies with instrumented data flows through seven distinct channels: final outputs, inter-agent messages, tool inputs/outputs, shared memory, logs, and artifacts.
@@ II-B Privacy Benchmarks and Evaluation Tools
Most recently, Green et al. [17] (EMNLP 2025) demonstrated that Large Reasoning Models frequently leak sensitive user data within their internal reasoning traces, even when the final output appears safe, a “leaky thoughts” phenomenon whereby increased test-time compute amplifies privacy risk. Their work establishes that output-only auditing is insufficient even for single-model deployments. AgentLeak extends this observation from the single-model reasoning level to the multi-agent architectural level: whereas Green et al. focus on the internal thinking of an individual model, we audit the internal coordination (C2 channels) between multiple agents, revealing a previously unquantified attack surface where inter-agent messages leak at $2.1\times$ the rate of final outputs.
@@ II-C Defense Mechanisms
Guardrail Systems. NVIDIA’s NeMo Guardrails [20] and Meta’s LlamaGuard [21] enforce policies at system interfaces but do not monitor internal communication pathways.
Privacy-Focused Prompts. Drawing on Constitutional AI [22] and privacy-preserving agent designs [23], these encode privacy norms in natural language. The hope is that agents internalize them. In practice, agents often ignore prompt instructions when coordinating with each other.
@@
TABLE I: Benchmark Landscape: Identifying the Evaluation Gap
Tool Multi-Agent Output Ch. Internal Ch. Type
@@
NeMo Guard. Defense
@@ III-B The Seven Leakage Channels
External Channels: Channel C1 captures final outputs presented to users, the primary target of current safety mechanisms. Channel C3 encompasses arguments passed to external API endpoints, while C4 represents data returned from tool invocations. Channel C6 covers telemetry streams and system logs. Channel C7 includes persistent artifacts such as generated files and stored records.
Internal Channels: Channel C2 comprises inter-agent messages exchanged during task delegation and coordination. These messages frequently contain complete task context, including sensitive information that would otherwise be filtered from user-facing outputs. Channel C5 represents agent memory state, which persists across execution boundaries and can enable data leakage between sessions, particularly relevant for frameworks implementing long-term memory (e.g., AutoGPT).
@@ III-C Privacy Leakage Formalization
Formally, let $\mathcal{V}=\{v_{1},\ldots,v_{n}\}$ be the private vault of $n$ sensitive fields we must protect. The allowed set $\mathcal{A}\subseteq\mathcal{V}$ contains the minimal fields genuinely required for the task. Let $\mathcal{C}=\{C_{1},\ldots,C_{7}\}$ denote our seven channels. We define a semantic similarity function $\operatorname{sim}(v,c)$ measuring the usage of vault field $v$ in channel content $c$. A leakage event occurs when:
$\exists\,v_{i}\in(\mathcal{V}\setminus\mathcal{A}),\,\exists\,C_{j}\in\mathcal{C}:\operatorname{sim}(v_{i},\operatorname{content}(C_{j}))>\tau\,$, (1)
@@
This turns data minimization [8], an abstract legal principle from GDPR [25], HIPAA [24], and Quebec’s Law 25 [26], into a testable empirical criterion.
TABLE II: Operational Definitions for Leakage and Risk Metrics
Concept Definition
Leakage vs. Exposure: presence of sensitive data in a channel
Exposure (C2/C5). Leakage: unauthorized exposure. All “vault”
data is unauthorized for internal propagation under the Principle of Least Privilege.
Fair Comparison When comparing single- vs. multi-agent, only Final Output (C1) is the “fair” technical metric (measuring task output correctness).
System Risk The “true” risk metric includes C2+C5, as these internal states are logged, stored, and potentially accessible to infrastructure providers or attackers (lateral movement), representing a regulatory liability (GDPR Art. 25).
@@ IV AgentLeak Benchmark Design
Figure 1: AgentLeak framework-agnostic evaluation harness. The Adapter Layer hooks into any multi-agent framework (LangChain, CrewAI, AutoGPT, MetaGPT, or custom stacks), normalizes trace events into a unified JSONL store covering all seven channels (C1–C7), and feeds them through a three-tier leakage detection pipeline (Presidio, LLM-as-Judge, Hybrid) to produce standardized privacy metrics (ELR, WLS, CLR, ASR).
Figure 2: AgentLeak’s multi-agent system architecture showing the seven leakage channels. Blue regions denote core pipeline components; the red-highlighted dashed box (C2) and red labels (C5) mark the primary internal leakage channels. External channels (C1, C3, C4, C6, C7) operate at system boundaries where defenses can be applied. Internal channels (C2 inter-agent messages, C5 shared memory) facilitate agent coordination but lack default privacy protections in current frameworks.
@@ IV-A Design Principles
Framework Independence. Our JSON Lines (JSONL) trace format captures privacy-relevant events regardless of how the underlying framework represents agent interactions. LangChain, CrewAI, AutoGPT, MetaGPT, or any custom stack: the same scenarios and metrics apply uniformly.
@@ IV-B Scenario Schema and Structure
The private vault lists every sensitive field that must be protected: social security numbers, diagnoses, account numbers, salaries, case details. Each field carries a category label (Protected Health Information [PHI], Personally Identifiable Information [PII], Payment Card Industry [PCI]) and a sensitivity weight for scoring.
@@
Ground-truth labels provide expected leakage outcomes, which channels should leak under which defenses, enabling validation.
@@ IV-C Domain Coverage
Our benchmark encompasses four industry verticals, 250 scenarios per domain, selected for their stringent privacy requirements and documented consequences of data breaches.
@@
Finance (250 scenarios): Financial scenarios span know-your-customer and anti-money laundering workflows, portfolio management, transaction disputes, credit underwriting, and regulatory reporting. Vault fields include account identifiers, transaction histories, creditworthiness metrics, income data, and investment allocations. Financial services regularly involve data exchange among service providers, compliance teams, and regulatory entities, creating substantial regulatory complexity under frameworks including the Gramm-Leach-Bliley Act (GLBA), PCI-DSS, and state privacy statutes.
Legal (250 scenarios): Legal practice scenarios address contract analysis, discovery document management, client onboarding, privilege protection, and case documentation. Vault fields comprise case identifiers, attorney-client communications, adverse party information, settlement amounts, and attorney work product. Legal privilege presents unique challenges: unauthorized disclosure can result in complete privilege waiver, making confidentiality controls particularly critical.
@@
Table III presents the complete scenario distribution across topology and threat level dimensions.
TABLE III: AgentLeak Scenario Distribution by Topology and Threat Level
Dimension Category Count Pct.
@@
Multi-agent (Coord.-Worker) 600 60%
@@ IV-D Data Generation and Validation
Explicit canaries: Unique markers, trivially detectable if leaked verbatim, baseline validation.
Format-valid fakes: Syntactically correct identifiers satisfying checksums (Luhn, SSN area codes), testing format-based detection.
@@
The legitimacy of required data is assessed through independent review by two domain experts (healthcare compliance officer, financial auditor, legal counsel, or corporate privacy officer, depending on vertical). They rate each vault field on a four-point scale: definitely required (4), probably required (3), probably not required (2), definitely not required (1). Fields rated 3+ by both annotators go into the allowed set. Inter-rater reliability is measured via Cohen’s $\kappa$ [28]: Healthcare 0.82, Finance 0.79, Legal 0.81, Corporate 0.75, all indicating substantial agreement ($\kappa>0.6$) across verticals.
@@ IV-E Three-Tier Detection Pipeline
Tier 1 (Canary Matching): Regex search for unique markers inserted as explicit canaries. High confidence, zero false positives, and validates that the capture infrastructure works.
@@ V Attack Taxonomy
Our 32-class taxonomy provides a standardized vocabulary for reproducible privacy research. We first motivate its necessity, then describe each attack family. Table IV outlines the mapping of attack families to the 32-class taxonomy.
@@ V-A Motivation
Prior work on data extraction [29], PII leakage analysis [30], and red teaming [31] targets static LLMs. Agentic systems introduce additional attack surfaces (tools, memory, and multi-agent coordination) that existing taxonomies do not cover [32]. A systematic taxonomy enables:
@@ V-B Attack Families
TABLE IV: Attack taxonomy: six families mapping to 32 classes
@@
F3: Memory and Persistence Attacks (5 classes) target the temporal aspect of how agents retain state. Once sensitive data lands in persistent storage, an attacker can extract it long after the original task has completed [37].
@@
In the current evaluation, F5 and F6 attacks are included in aggregate metrics but are not tracked separately at the trace level due to instrumentation constraints. We report per-family Attack Success Rate (ASR) for F1 through F4 only. Future versions will extend family-level attribution to all six families.
Table V maps attack families to adversary capability levels, showing which attacks are available at each threat level.
TABLE V: Attack Family to Adversary Level Mapping
@@ VI-B Detection Pipeline Validation
To mitigate evaluation leakage, judges only receive minimally necessary context: a trace segment and a masked scenario identifier. Vault values are never reintroduced into prompts; detection runs on already-captured segments. Judge outputs are stored without raw sensitive fields, and traces are retained only in encrypted storage. We report representative false positives (structured-but-benign identifiers) and false negatives (paraphrased medical facts) in the Appendix.
@@ VI-D Experimental Setup
The benchmark comprises 1,000 designed scenarios (250 per vertical: healthcare, finance, legal, corporate). Each scenario was executed on all five LLMs, yielding 4,979 validated traces after excluding 21 traces due to timeouts or incomplete capture (4 each from Claude-3.5-Sonnet, GPT-4o, Llama-3.3-70B, and Mistral-Large; 5 from GPT-4o-mini), resulting in a total of n=4,979n=4{,}979 model–scenario pairs. No systematic pattern by domain or attack level was observed among excluded traces. Table VI summarizes the trace distribution.
TABLE VI: Benchmark Execution Summary
@@
Scenarios/vertical 250 GPT-4o 996 Finance 1,244
@@
All evaluations used temperature =0.7=0.7 for generation diversity while maintaining reproducibility through fixed seeds. Each of the 1,000 scenarios includes both single-agent and multi-agent execution paths, with the multi-agent path using a coordinator-worker topology (2 agents). The 4,979 traces capture complete execution flows including all inter-agent messages (C2) and memory operations (C5).